# ICAGC 2024: Inspirational and Convincing Audio Generation Challenge 2024

### v1

@@ Abstract
The Inspirational and Convincing Audio Generation Challenge 2024 (ICAGC 2024) is part of the ISCSLP 2024 Competitions and Challenges track [1]. While current text-to-speech (TTS) technology can generate high-quality audio, its ability to convey complex emotions and controlled detail content remains limited. This constraint leads to a discrepancy between the generated audio and human subjective perception in practical applications like companion robots for children and marketing bots. The core issue lies in the inconsistency between high-quality audio generation and the ultimate human subjective experience. Therefore, this challenge aims to enhance the persuasiveness and acceptability of synthesized audio, focusing on human alignment convincing and inspirational audio generation.
@@ 1 Introduction
The Inspirational and Convincing Audio Generation Challenge 2024 (ICAGC 2024) is a pioneering event in the realm of synthetic audio, set to be a highlight within the ISCSLP 2024 Competitions and Challenges track. Despite the significant advancements in text-to-speech (TTS) technology, current systems often fall short in conveying intricate emotions and controlled, detailed content, which is crucial for applications such as companion robots for children and marketing bots. This gap leads to a noticeable discrepancy between the generated audio and human subjective perception, particularly in real-world applications.
@@ 5.5 Notice
6 Registration
If you are registering as a team, it is sufficient for one team member to fill the registration Google form, they will be contacted for all communications.
https://forms.gle/vgpWTZMK8noacC3GA
@@ 7.2 Submission Procedure
8 Challenge Ranks and Rewards
  • Ranks:Track 1 and Track 2 are ranked separately.
  • Certificate: Top 5 ranked teams from each Track will receive certificates.
  • Souvenir:Top 3 ranked teams from each Track will gain exquisite gifts.
  • Paper: Taking into account the challenge ranks and submitted papers comprehensively, 5 teams will be invited to present their work at ISCSLP 2024.
  • Internship or referral opportunities
9 Important Dates
  • May 15, 2024: Registration for the challenge opens
  • May 25, 2024: Release voice cloning reference speaker data
  • June 20, 2024: Release text test data
  • June 30, 2024: Submission deadline
  • July 10, 2024: Announcement of winners in all tracks
  • July 20, 2024: Grand challenge 5-page papers due (by invitation only - challenge organizers will invite the top 5 ranked teams)
  • July 30, 2024: Camera-ready paper upload dead
  • Aug 15, 2024: Author’s registration deadline
Note: The dates given apply to any time zone across the world.

### v2

@@ Abstract
The Inspirational and Convincing Audio Generation Challenge 2024 (ICAGC 2024) is part of the ISCSLP 2024 Competitions and Challenges track. While current text-to-speech (TTS) technology can generate high-quality audio, its ability to convey complex emotions and controlled detail content remains limited. This constraint leads to a discrepancy between the generated audio and human subjective perception in practical applications like companion robots for children and marketing bots. The core issue lies in the inconsistency between high-quality audio generation and the ultimate human subjective experience. Therefore, this challenge aims to enhance the persuasiveness and acceptability of synthesized audio, focusing on human alignment convincing and inspirational audio generation. A total of 19 teams have registered for the challenge, and the results of the competition and the competition are described in this paper.
@@ 1 Introduction
The Inspirational and Convincing Audio Generation Challenge 2024 (ICAGC 2024) is a pioneering event in the realm of synthetic audio, set to be a highlight within the ISCSLP 2024 Competitions and Challenges track [1]. Despite the significant advancements in text-to-speech (TTS) technology, current systems often fall short in conveying intricate emotions and controlled, detailed content, which is crucial for applications such as companion robots for children and marketing bots. This gap leads to a noticeable discrepancy between the generated audio and human subjective perception, particularly in real-world applications.
@@ 6.2 Submission Procedure
Table 1: Track 1 Result
Team Name Quality Similarity Emotion MOS(avg) MOS(std)
{CJK}UTF8gbsn油菜花队伍 3.89 3.83 3.85 3.86 0.22
NPU-HWC 3.84 3.48 3.58 3.63 0.30
zyzx_ai 3.67 3.50 3.50 3.56 0.42
PeiyangTTSer 3.64 3.52 3.28 3.48 0.39
Happy Happy 3.33 3.45 3.33 3.37 0.36
Table 2: Track 2 Result
Team Name Similarity Emotion Matching MOS(avg) MOS(std)
NPU-HWC 4.15 3.66 3.33 3.67 0.68
PeiyangTTSer 3.33 3.33 2.66 3.06 0.41
7 Results
7.1 Summary of submitted systems and evaluation results for Track 1
Track 1 aims to clone a given speaker’s voice and adjust the synthesized speech’s emotion according to different topics. Team 1 introduces an expressive TTS synthesizer to address the Track 1 challenge. The entire framework is a cascaded system. First, the LLM model predicts speech tokens constructed by multilingual Wav2Vec and K-means. Then, a variational autoencoder (VAE) as a voice conversion model is applied to generate the mel spectrogram from the speech tokens. Finally, the HifiGAN vocoder is used to synthesize the waveform using the generated mel spectrogram as input. To encourage the disentanglement of timbre and style, their strategy is to limit the responsibilities of the autoregressive model to capturing phoneme content, duration, and prosody, while designing a separate token-to-mel decoder for reconstructing the speaker’s identity. Additionally, they propose a data processing pipeline to transform in-the-wild speech data into high-quality training data with annotations for speech generation. Team 2 implements a zero-shot speaker and style cloning speech generator in Track 1, similar to Tortoise TTS, specifically designed for zero-shot speaker and style cloning. They use Single-Codec as the speech tokenizer for their proposed method, which explicitly decouples timbre and speaking style at the token level, thereby reducing the acoustic modeling burden on the autoregressive model. Additionally, they adopt DSPGAN to upsample 16k mel spectrograms to 48k high-fidelity waveforms. Team 3 mainly uses the pre-trained GPT-SoVITS, a large-scale, two-stage speech synthesis system. The system structure includes two parts: an autoregressive (AR) model and a non-autoregressive (NAR) model. The AR model focuses on extracting semantic information from the synthesized speech, such as content and style, while the NAR model focuses on controlling non-linguistic elements in the synthesized output, such as timbre and emotion. They propose a few-shot learning strategy for specific topics, fine-tuning the pre-trained model with sentences from different speakers on the same topic to optimize the model for particular themes and individual speaker characteristics. Team 4 proposes Hola-TTS, a robust speech synthesis model for cross-lingual zero-shot speech synthesis, suitable for four languages: Chinese, English, Japanese, and Korean. Hola-TTS uses a training approach based on large language models (LLMs). It first discretizes the speech signal into discrete tokens using the EnCodec encoder. Simultaneously, text input is converted into a sequence of phonemes with phoneme-level language labels. Both are concatenated and fed into a Transformer decoder for predicting the next token. During inference, tokens are predicted in an autoregressive manner, and the EnCodec decoder reconstructs the audio. Specifically, while large models are not particularly sensitive to data quality, clean data still improves the quality of the generated output. Therefore, they perform data cleaning before training. The result of Track1 are listed in Table. 1.
7.2 Summary of submitted systems and evaluation resultsfor Track 2
Track 2 aims to add realistic background sounds to the speech, making the audio more lifelike and convincing in real-life scenarios. Team 5 develops a large language model (LLM)-based cascading system to generate appropriate background audio for speech synthesis, consisting mainly of two stages: text-to-description and description-to-audio. In the text-to-description stage, the goal is to use the LLM’s semantic understanding to generate descriptive text corresponding to the speech transcript. For this, they select DeepSeek-V22, a robust expert mixture model (MoE) known for its strong language understanding capabilities. In the description-to-audio stage, they use Tango 2 to generate the background audio. Additionally, since the main focus of the audio is the speech content, they apply a 10dB gain reduction to the generated background audio during the mixing process to ensure it does not overshadow the primary speech. Team 6 builds an innovative system that deeply integrates the GPT-SoVITS1 and AudioLDM 2 models to achieve highly emotional and realistic background audio for text-to-speech synthesis. They innovatively combine GPT-3.5 with AudioLDM 2 in Track 2. Based on the results achieved from fine-tuning GPT-SoVITS, the system uses LLM’s deep understanding of text content to intelligently match each segment of speech with the most appropriate background description. Then, AudioLDM 2 generates these background sounds, ensuring they perfectly blend with the speech in terms of style, rhythm, and other aspects. The result of Track2 are listed in Table. 2.
8 Conclusion
The ICAGC 2024 Challenge successfully advances the development of synthetic audio technology, particularly in enhancing the emotional conveyance and persuasiveness of audio. Through the two tracks, we observe the application of various innovative approaches that not only improve the quality of synthetic audio but also enhance emotional resonance with listeners. The diversity and innovation of the entries indicate that TTS technology is evolving towards a more humanized and personalized direction. The results of the challenge highlight the important role of evaluation metrics in driving technological development.